# Hidden monotonicity and canonical transformations for mean field games and master equations

### v1

@@ 1. Introduction
(1.1) $\left\{\begin{array}[]{rl}-\partial_{t}V(t,x,\mu)+H(x,\mu,\partial_{x}V)-\mathcal{N}V(t,x,\mu)-\frac{\beta^{2}}{2}\Delta_{{\rm ind}}V-\frac{\beta_{0}^{2}}{2}\Delta_{{\rm com}}V(t,x,\mu)&=0,\\[3.0pt] &{\rm{in}}\ (0,T)\times\mathbb{R}^{d}\times{\mathscr{P}}_{2}(\mathbb{R}^{d}),\\[3.0pt] V(T,x,\mu)&=G(x,\mu),\\[3.0pt] &{\rm{in}}\ \mathbb{R}^{d}\times{\mathscr{P}}_{2}(\mathbb{R}^{d}),\end{array}\right.$
@@
$\mathcal{N}V(t,x,\mu)$ $=-\int_{\mathbb{R}^{d}}\partial_{\mu}V(t,x,\mu,\tilde{x})\cdot\partial_{p}H(\tilde{x},\mu,\partial_{x}V(t,\tilde{x},\mu)){\mathrm{d}}\mu(\tilde{x})$
$\Delta_{{\rm ind}}V$ $=\operatorname{tr}(\partial_{xx}V(t,x,\mu))+\int_{\mathbb{R}^{d}}\operatorname{tr}(\partial_{\tilde{x}\mu}V(t,x,\mu,\tilde{x})){\mathrm{d}}\mu(\tilde{x})$
@@
$\Delta_{{\rm com}}V$ $=\operatorname{tr}(\partial_{xx}V(t,x,\mu))+\int_{\mathbb{R}^{d}}\operatorname{tr}(\partial_{\tilde{x}\mu}V(t,x,\mu,\tilde{x})){\mathrm{d}}\mu(\tilde{x})+2\int_{\mathbb{R}^{d}}\operatorname{tr}(\partial_{x\mu}V(t,x,\mu,\tilde{x})){\mathrm{d}}\mu(\tilde{x})$
@@
In [CD18b, Theorem 5.46] the authors have shown that the master equation (1.1) is globally well-posed if the data are LL monotone and possess additional regularity assumptions. Several other works provide similar conclusions. We refer to [CDLL19, Theorem 2.4.5] for the case when the physical space is the flat torus instead of $\mathbb{R}^{d}$ and to [CCD22, Theorems 56 and 58] to the case without common noise (i.e. $\beta_{0}=0$). We refer also to [JR23] for new results and clarifications regarding the results from [CDLL19]. However, [CD18b, Theorem 5.46] is the closest result for our purposes.
@@
In the context of displacement monotonicity global in time well-posedness have been obtained chronologically as follows. [GM22a] provided this in the context of deterministic and potential (in particular $\beta=\beta_{0}=0$ and $H$ separable) games (for similar results, see also [BGY24]). [GMMZ22] provided the first global in time well-posedness result in the case of non-separable displacement monotone Hamiltonians and non-degenerate idiosyncratic noise (i.e. $\beta\neq 0$). Finally, [BMM23] provided the result in the case of degenerate idiosyncratic noise (i.e. $\beta=0$) and compared to [GMMZ22], under lower level regularity assumptions on the data, and the weaker version of the displacement monotonicity condition on $H$.
@@
Several other recent developments have seen the light in the context of the well-posedness of MFG master equations. For a non-exhaustive list we refer to [AM23, Ber21, CCP23, CD24, GM22b, GM23].
Our contributions. In this paper our main objective is to present new global well-posedness theories for the master equation (1.1). The heart of our analysis consist of so-called canonical transformations which in particular reveal new perspectives on existing and new monotonicity conditions on the Hamiltonians and final data associated to (1.1), and in turn lead to new well-posedness theories. The values of the noise intensities, $\beta,\beta_{0}$ will not not be significant in our consideration, and our main results hold true also for degenerate problems, i.e. when $\beta=0$ or $\beta_{0}=0$.
In classical Hamiltonian mechanics, canonical transformations are coordinate transformations on the phase space, which preserve the structure of Hamilton’s equations. In symplectic geometry, canonical transforms are known as symplectomorphisms (where the phase space is a cotangent bundle and the symplectic form is the canonical 2-form). Since in our setting we are only concerned with Euclidean space we do not use the symplectic terminology. However, one could use symplectomorphisms to generate new well-posedness theories for Hamilton–Jacobi equations and the master equation in more general settings (i.e. when the underlying space is not Euclidean). We refer the reader to [Arn89] for a introduction to applications of symplectic geometry in classical mechanics. We refer also to our companion short note [BM24], where we explain the regularization effect of such transformations in the case of deterministic finite dimensional HJB equations.
As the master equation has in particular a natural character arising from infinite dimensional Hamiltonian dynamics, we will show below, that such transformations play a deep role in obtaining new well-posedness theories for it.
@@
This is a ‘finite dimensional’ transformation, as there is no change in the measure variable $\mu$. Having defined these transformations, we can formulate our first main result.
@@
The message of this theorem is that if one produces a well-posedness theory for the master equation, this will lead to a whole one parameter family of well-posedness theories, with the transformed data. A deeper consequence of this theorem is the opposite implication. Suppose that one is given the data $(H,G)$. If one is able to find a suitable range of the parameter $\alpha$ such that $(H_{\alpha},G_{\alpha})$ satisfies some well-known monotonicity conditions, then the problem with the original data must be well-posed. This second one will be the direction that we investigate in this paper.
Fix $\alpha\in\mathbb{R}$. It is easy to see that $G$ is LL monotone, if and only if $G_{\alpha}$ is LL monotone and the situation is the same for separable $H$. However, as we will show below, this phenomenon is much different in the displacement monotone regime. Therefore Theorem 1.1 has powerful applications in the context of displacement monotonicity but not for LL monotonicity.
In Theorem 2.9, we propose easily verifiable sufficient conditions on $H$ to ensure that $H_{\alpha}$ is displacement monotone.
This theorem has an immediate consequence which can informally be formulated as follows. This result provides a new global well-posedness theory for the master equation.
@@
Corollary 1.2 has a deep message: if the Hamiltonian is such that $\partial_{xp}H$ is sufficiently large compared to other second order derivatives of $H$ and the second order derivatives of $G$, then we have a global well-posedness theory for the master equation. Therefore $\partial_{xp}H$, and in particular adding suitable multiples of the function $(x,p,\mu)\mapsto p\cdot x$ to $H$ have a ‘regularization effect’. By carefully examining Lemma 2.7, we see that what is going on is that the $p\cdot x$ term is transformed into a multiple of $\frac{\left|x\right|^{2}}{2}$, which provides displacement monotonicity for the problem and hence regularizes the master equation. It is easy to see that adding a suitable multiple of the term $\frac{\left|x\right|^{2}}{2}$ to $H$ produces displacement monotonicity. Clearly, these regularization effects are independent of the noise intensities.
@@
In the recent paper [MZ24], the authors proposed a notion of anti-monotonicity for final data $G$. They have described some heavy sufficient conditions on $H$ and $G$ which result in a global well-posedness theory of (1.1), if $\beta\neq 0$, and $G$ is suitably anti-monotone. There was an emphasis on the fact that $G$ needed to be ‘sufficiently’ anti-monotone.
It turns out that these well-posedness results from [MZ24], under the additional assumptions that $H$ is strictly convex in the $p$-variable fall directly into the framework of the canonical transformations and they are an easy consequence of our main results, from Corollary 1.2. More precisely, first in Proposition 2.11 we show that if $G$ is $\lambda$-anti-monotone, this implies that it is displacement semi-monotone with a constant which depends only on $\lambda$ (in particular, the displacement semi-monotonicity constant is independent of the second derivative bounds of $G$). Having strong convexity of $H$ in the $p$-variable, which has also bounded second derivatives allows us to use our Corollary 1.2. The Hamiltonian considered in [MZ24] has the form of
@@
for some constant matrix $A_{0}\in\mathbb{R}^{d\times d}$. This is slightly different than $\tilde{H}$ from our Corollary 1.2, but the term $\langle A_{0}p,x\rangle$ has exactly the same role as $\alpha p\cdot x$ in our consideration. Therefore, for completeness, as our last contributions, in Proposition 2.15 and Remark 2.16 we show that the assumptions from the main theorem in [MZ24] essentially imply our assumptions. Furthermore, in the case of Hamiltonians which are strongly convex in the $p$-variable, our results need less and weaker assumption, and they hold true without the presence of a non-degenerate idiosyncratic noise. In particular, we demonstrate that the emphasis on the sufficient anti-monotonicity of $G$ in [MZ24] is misleading, and this is not needed. Specifically, in [MZ24] it is remarked: “…we will need to require our data to be sufficiently anti-monotone in appropriate sense”. However we will see that anti-monotonicty is not needed (as anti-monotonicity implies semi-monotonicity) and that [MZ24] has other other, more essential assumptions on $H$ which are what really give the well-posedness result.
@@
    $H_{\varphi}(x,\mu,p):=H(x,\mu,x-\nabla\varphi(x))+\frac{\beta^{2}+\beta_{0}^{2}}{2}\Delta\varphi(x)\ \ {\rm{and}}\ \ G_{\varphi}(x,\mu):=G(x,\mu)+\varphi(x).$
@@
    We remark that in general Hamiltonians are only defined up to an additive constant. In classical mechanics, this is saying that we may pick any value to correspond to the ‘zero energy’. In the presence of noise the attentive reader will notice that our $H_{\alpha}$ is not the same as the $H_{\varphi}$ defined above, when $\varphi(x)$ is taken to be $\frac{\alpha}{2}|x|^{2}$ However, this is not an issue as the difference between the two is a constant. In particular, the two Hamiltonians are equivalent. Thus, we could have defined our $H_{\alpha}$ as $H_{\alpha}(x,\mu,p):=H(x,\mu,p-\alpha x)+{\frac{(\beta^{2}+\beta_{0}^{2})d}{2}}\alpha$ which would then be the exact same as $H_{\varphi}$ defined above, however this would introduce unnecessary notational clutter.
@@
2. Preliminaries and new well-posedness theories for master equations
2.1. Some notations
@@
    $\int_{\mathbb{R}^{d}}\langle\partial_{xx}G(x,\mu)\xi(x),\xi(x)\rangle{\mathrm{d}}\mu(x)+\int_{\mathbb{R}^{d}\times\mathbb{R}^{d}}\langle\partial_{x\mu}G(x,\mu,\tilde{x})\xi(x),\xi(\tilde{x})\rangle{\mathrm{d}}\mu(x){\mathrm{d}}\mu(\tilde{x})\geq 0,$
@@
    (2.2) $\int_{\mathbb{R}^{d}\times\mathbb{R}^{d}}\left[\partial_{x\mu}H(x,\mu,\tilde{x},p(x))v(\tilde{x})+\partial_{xx}H(x,\mu,p(x))v(x)\right]\cdot v(x){\mathrm{d}}\mu(x){\mathrm{d}}\mu(\tilde{x})$
    $+\frac{1}{4}\int_{\mathbb{R}^{d}}\left\{\Big{|}[\partial_{pp}H(x,\mu,p(x))]^{-\frac{1}{2}}\int_{\mathbb{R}^{d}}\partial_{p\mu}H(x,\mu,\tilde{x},p(x))v(\tilde{x}){\mathrm{d}}\mu(\tilde{x})\Big{|}^{2}\right\}{\mathrm{d}}\mu(x)\leq 0,$
@@
$\lambda_{0}\int_{\mathbb{R}^{d}}\langle\partial_{xx}G(x,\mu)\xi(x),\xi(x)\rangle{\mathrm{d}}\mu(x)+\lambda_{1}\int_{\mathbb{R}^{d}\times\mathbb{R}^{d}}\langle\partial_{x\mu}G(x,\mu,\tilde{x})\xi(x),\xi(\tilde{x})\rangle{\mathrm{d}}\mu(x){\mathrm{d}}\mu(\tilde{x})$
@@
2.2. New Well-Posedness Theories for MFG and master equations
@@
    Assumptions 1 and 2 from above are the standing assumptions imposed in [BMM23].
We now prove Theorem 1.1.
Proof of Theorem 1.1.
@@
$+\int_{\mathbb{R}^{d}\times\mathbb{R}^{d}}\left[\Big{(}\partial_{x\mu}H(x,\mu,\tilde{x},p(x))-2\alpha\partial_{p\mu}H(x,\mu,\tilde{x},p(x))\Big{)}v(\tilde{x})\right]\cdot v(x){\mathrm{d}}\mu(x){\mathrm{d}}\mu(\tilde{x})$
$+\frac{1}{4}\int_{\mathbb{R}^{d}}\left\{\bigg{|}[\partial_{pp}H(x,\mu,p(x))]^{-\frac{1}{2}}\left[\int_{\mathbb{R}^{d}}\partial_{p\mu}H(x,\mu,\tilde{x},p(x))v(\tilde{x}){\mathrm{d}}\mu(\tilde{x})+2\alpha\partial_{pp}H(x,\mu,p(x))v(x)\right]\bigg{|}^{2}\right\}{\mathrm{d}}\mu(x)$
@@
$\partial_{xx}\tilde{H}(x,\mu,p)$ $=\partial_{xx}H(x,\mu,p-\alpha x)-2\alpha\operatorname{Re}(\partial_{xp}H(x,\mu,p-\alpha x))+\alpha^{2}\partial_{pp}H(x,\mu,p-\alpha x),$
@@
(2.4) $\int_{\mathbb{R}^{d}\times\mathbb{R}^{d}}\left[\partial_{x\mu}H(x,\mu,\tilde{x},p(x))v(\tilde{x})-\alpha\partial_{p\mu}H(x,\mu,\tilde{x},p(x))v(\tilde{x})\right]\cdot v(x){\mathrm{d}}\mu(x){\mathrm{d}}\mu(\tilde{x})$
$+\int_{\mathbb{R}^{d}}\left[\partial_{xx}H(x,\mu,p(x))v(x)-2\alpha\partial_{xp}H(x,\mu,p(x))v(x)+\alpha^{2}\partial_{pp}H(x,\mu,p(x))v(x)\right]\cdot v(x){\mathrm{d}}\mu(x)$
$+\frac{1}{4}\int_{\mathbb{R}^{d}}\left\{\Big{|}[\partial_{pp}H(x,\mu,p(x))]^{-\frac{1}{2}}\int_{\mathbb{R}^{d}}\partial_{p\mu}H(x,\mu,\tilde{x},p(x))v(\tilde{x}){\mathrm{d}}\mu(\tilde{x})\Big{|}^{2}\right\}{\mathrm{d}}\mu(x)\leq 0,$
@@
$\left|\partial_{x\mu}H\right|:=\sup_{(x,\mu,p,\tilde{x})\in\mathbb{R}^{d}\times{\mathscr{P}}_{2}(\mathbb{R}^{d})\times\mathbb{R}^{d}\times\mathbb{R}^{d}}\left|\partial_{x\mu}H(x,\mu,p,\tilde{x})\right|;$
$\left|\partial_{p\mu}H\right|:=\sup_{(x,\mu,p,\tilde{x})\in\mathbb{R}^{d}\times{\mathscr{P}}_{2}(\mathbb{R}^{d})\times\mathbb{R}^{d}\times\mathbb{R}^{d}}\left|\partial_{p\mu}H(x,\mu,p,\tilde{x})\right|;$
@@
Suppose that $H:\mathbb{R}^{d}\times{\mathscr{P}}_{2}(\mathbb{R}^{d})\times\mathbb{R}^{d}\to\mathbb{R}$ satisfies $\partial_{pp}H(x,\mu,p)\geq c_{0}^{-1}I$, for some $c_{0}>0$ and for all $(x,\mu,p)\in\mathbb{R}^{d}\times{\mathscr{P}}_{2}(\mathbb{R}^{d})\times\mathbb{R}^{d}$. Suppose that $\underline{\kappa}(\partial_{xp}H),\left|\partial_{pp}H\right|,\left|\partial_{xx}H\right|,\left|\partial_{p\mu}H\right|$ and $\left|\partial_{x\mu}H\right|$ are finite. Define
@@
$\int_{\mathbb{R}^{d}\times\mathbb{R}^{d}}\left[\partial_{x\mu}H(x,\mu,\tilde{x},p(x))v(\tilde{x})-\alpha\partial_{p\mu}H(x,\mu,\tilde{x},p(x))v(\tilde{x})\right]\cdot v(x){\mathrm{d}}\mu(x){\mathrm{d}}\mu(\tilde{x})$
$+\int_{\mathbb{R}^{d}}\left[\partial_{xx}H(x,\mu,p(x))v(x)-2\alpha\partial_{xp}H(x,\mu,p(x))v(x)+\alpha^{2}\partial_{pp}H(x,\mu,p(x))v(x)\right]\cdot v(x){\mathrm{d}}\mu(x)$
@@
$\leq\int_{\mathbb{R}^{d}\times\mathbb{R}^{d}}[\left|\partial_{x\mu}H\right|+\alpha\left|\partial_{p\mu}H\right|+\left|\partial_{xx}H\right|-2\alpha\underline{\kappa}(\partial_{xp}H)+\alpha^{2}\left|\partial_{pp}H\right|]{\mathrm{d}}\mu(x){\mathrm{d}}\mu(\tilde{x})$
@@
2.2.1. Our results and previous results on the master equation involving displacement semi-monotone data

### v2

@@ 1. Introduction
(1.1) $\left\{\begin{array}[]{rl}-\partial_{t}V(t,x,\mu)+H(x,\mu,\partial_{x}V)+\mathcal{N}V(t,x,\mu)\\ -\frac{\beta^{2}}{2}\Delta_{{\rm ind}}V-\frac{\beta_{0}^{2}}{2}\Delta_{{\rm com}}V(t,x,\mu)&=0,\\[3.0pt] &{\rm{in}}\ (0,T)\times\mathbb{R}^{d}\times{\mathscr{P}}_{2}(\mathbb{R}^{d}),\\[3.0pt] V(T,x,\mu)&=G(x,\mu),\\[3.0pt] &{\rm{in}}\ \mathbb{R}^{d}\times{\mathscr{P}}_{2}(\mathbb{R}^{d}),\end{array}\right$.
@@
[…]
@@
[…]
@@
In [CD18b, Theorem 5.46] the authors have shown that the master equation (1.1) is globally well-posed if the data are LL monotone and possess additional regularity assumptions. Several other works provide similar conclusions. We refer to [CDLL19, Theorem 2.4.5] for the case when the physical space is the flat torus instead of $\mathbb{R}^{d}$ and to [CCD22, Theorems 56 and 58] to the case without common noise (i.e. $\beta_{0}=0$). We refer also to [JR25] for new results and clarifications regarding the results from [CDLL19]. However, [CD18b, Theorem 5.46] is the closest result for our purposes.
@@
In the context of displacement monotonicity global in time well-posedness have been obtained chronologically as follows. [GM22] provided this in the context of deterministic and potential (in particular $\beta=\beta_{0}=0$ and $H$ separable) games (for similar results, see also [BGY24]). [GMMZ22] provided the first global in time well-posedness result in the case of non-separable displacement monotone Hamiltonians and non-degenerate idiosyncratic noise (i.e. $\beta\neq 0$). Finally, [BMM25] provided the result in the case of degenerate idiosyncratic noise (i.e. $\beta=0$) and compared to [GMMZ22], under lower level regularity assumptions on the data, and the weaker version of the displacement monotonicity condition on $H$.
@@
Several other recent developments have seen the light in the context of the well-posedness of MFG master equations. For a non-exhaustive list we refer to [AM23, Ber21, CCP23, CD24, GM24, GM23].
Our contributions. In this paper our main objective is to explore some geometric features of Hamiltonian systems which could lead to the global well-posedness of the master equation (1.1). The heart of our analysis consist of so-called canonical transformations which in particular reveal new perspectives on existing and new monotonicity conditions on the Hamiltonians and final data associated to (1.1), and in turn lead to new well-posedness theories. The values of the noise intensities, $\beta,\beta_{0}$ will not not be significant in our consideration, and our main results hold true also for degenerate problems, i.e. when $\beta=0$ or $\beta_{0}=0$.
In classical Hamiltonian mechanics, canonical transformations are coordinate transformations on the phase space, which preserve the structure of Hamilton’s equations. In symplectic geometry, canonical transforms are known as symplectomorphisms (where the phase space is a cotangent bundle and the symplectic form is the canonical 2-form). Since in our setting we are only concerned with Euclidean space we do not use the symplectic terminology. However, it would be interesting to study how symplectomorphisms could potentially generate new well-posedness theories for Hamilton–Jacobi equations and the master equation in more general settings (i.e. when the underlying space is not Euclidean). We refer the reader to [Arn89] for a introduction to applications of symplectic geometry in classical mechanics. We refer also to our companion short note [BM25], where we explain the regularization effect of such transformations in the case of deterministic finite dimensional HJB equations.
As the master equation has in particular a natural character arising from infinite dimensional Hamiltonian dynamics, we will show below, that such transformations play a deep role in revealing hidden features of it.
@@
This is a ‘finite dimensional’ transformation, as there is no change in the measure variable $\mu$. Having defined these transformations, the heart of our analysis is based on the following observation: fix any $\alpha\in\mathbb{R}$, then the master equation with data $(H,G)$ is well-posed if and only if it is well-posed with data $(H_{\alpha},G_{\alpha})$ (see Theorem 3.2; in particular the solutions to the corresponding master equations differ only by an explicit function of $(t,x)$, parametrized by $\alpha$).
The message of this result is that if one produces a well-posedness theory for the master equation, this will lead to a whole one parameter family of well-posedness theories, with the transformed data. A deeper consequence of this theorem is the opposite implication. Suppose that one is given the data $(H,G)$. If one is able to find a suitable range of the parameter $\alpha$ such that $(H_{\alpha},G_{\alpha})$ satisfies some well-known monotonicity conditions, then the problem with the original data must be well-posed. This second one will be the direction that we investigate in this paper.
Fix $\alpha\in\mathbb{R}$. It is easy to see that $G$ is LL monotone, if and only if $G_{\alpha}$ is LL monotone and the situation is the same for separable $H$. However, as we will show below, this phenomenon is much different in the displacement monotone regime. Therefore the previously described result has powerful applications in the context of displacement monotonicity but not for LL monotonicity.
In the main theorem of this paper, Theorem 3.6, we propose easily verifiable sufficient conditions on $H$ to ensure that $H_{\alpha}$ is displacement monotone. As a consequence, we discover new regimes of global well-posedness of the master equation. In an informal way, this result can be summarized as follows (we refer to Theorem 3.6 for the precise statement).
@@
Suppose that $H:\mathbb{R}^{d}\times{\mathscr{P}}_{2}(\mathbb{R}^{d})\times\mathbb{R}^{d}\to\mathbb{R}$ is twice continuously differentiable with uniformly bounded second order derivatives. Suppose moreover that $H$ is strongly convex in the $p$-variable.
Suppose that the symmetric part of $\partial_{xp}H$ is bounded below by an explicit quantity depending on the other second derivatives of $H$. Then, $H_{\alpha}$ is displacement monotone for a suitable range of $\alpha\in\mathbb{R}$, depending on the size of the second derivatives of $H$ in a precise way.
Furthermore, if $G:\mathbb{R}^{d}\times{\mathscr{P}}_{2}(\mathbb{R}^{d})\to\mathbb{R}$ is twice continuously differentiable and displacement $\alpha$-monotone for such specific $\alpha$, then the master equation is globally well-posed.
This theorem has an immediate implication, coming from a sort of ‘regularization phenomenon’ of $\partial_{xp}H$. This can informally be formulated as follows.
@@
Corollary 1.2 has a deep message: if the Hamiltonian is such that $\partial_{xp}H$ is sufficiently large compared to other second order derivatives of $H$ and the second order derivatives of $G$, then we have a global well-posedness theory for the master equation. Therefore $\partial_{xp}H$, and in particular adding suitable multiples of the function $(x,p,\mu)\mapsto p\cdot x$ to $H$ can produce a ‘regularization effect’ for the master equation, independently of $T>0$. By carefully examining Lemma 3.4, we see that what is going on is that the $p\cdot x$ term is transformed into a multiple of $\frac{\left|x\right|^{2}}{2}$, which provides displacement monotonicity for the problem and hence regularizes the master equation. It is easy to see that adding a suitable multiple of the term $\frac{\left|x\right|^{2}}{2}$ to $H$ produces displacement monotonicity. Clearly, these regularization effects are independent of the noise intensities.
Remark 1.4.
We emphasize that the regularization provided by the function $(x,p,\mu)\mapsto\alpha p\cdot x$ in the statement of Corollary 1.2 produces indeed a genuinely new class of data, not covered in the literature before, for which the master equation is globally well-posed. In particular, if we take an arbitrary pair of data $(H,G)$, not satisfying any monotonicity condition (either displacement or LL, if $H$ is separable), it is immediate to check that $\tilde{H}$ will satisfy neither displacement monotonicity nor LL monotonicity. Therefore, the monotonicity of the pair $(\tilde{H},G)$ is indeed hidden.
@@
In the recent paper [MZ24], the authors proposed a notion of anti-monotonicity for final data $G$. They have described some sufficient conditions on $H$ and $G$ which result in a global well-posedness theory of (1.1), if $\beta\neq 0$, and $G$ is suitably anti-monotone. There was an emphasis on the fact that $G$ needed to be ‘sufficiently’ anti-monotone.
It turns out that these well-posedness results from [MZ24], under the additional assumptions that $H$ is strictly convex in the $p$-variable fall directly into the framework of the canonical transformations and they are an easy consequence of our main results, in particular Corollary 1.2. More precisely, first in Proposition 3.8 we show that if $G$ is $\lambda$-anti-monotone, this implies that it is displacement semi-monotone with a constant which depends only on $\lambda$ (in particular, the displacement semi-monotonicity constant is independent of the second derivative bounds of $G$). Having strong convexity of $H$ in the $p$-variable, which has also bounded second derivatives allows us to use our Corollary 1.2. The Hamiltonian considered in [MZ24] has the form of
@@
for some constant matrix $A_{0}\in\mathbb{R}^{d\times d}$. This is slightly different than $\tilde{H}$ from our Corollary 1.2, but the term $\langle A_{0}p,x\rangle$ has exactly the same role as $\alpha p\cdot x$ in our consideration. Therefore, for completeness, as our last contributions, in Proposition 3.12 and Remark 3.13 we show that the assumptions from the main theorem in [MZ24] essentially imply our assumptions. Furthermore, in the case of Hamiltonians which are strongly convex in the $p$-variable, our results need less and weaker assumption, and they hold true without the presence of a non-degenerate idiosyncratic noise. In particular, we demonstrate that the emphasis on the sufficient anti-monotonicity of $G$ in [MZ24] is misleading, and this is not needed. Specifically, in [MZ24] it is remarked: “…we will need to require our data to be sufficiently anti-monotone in appropriate sense”. However we will see that anti-monotonicty is not needed (as anti-monotonicity implies semi-monotonicity) and that [MZ24] has other, more essential assumptions on $H$ which are what really give the well-posedness result.
We would like to emphasize that in this paper we provide a general mechanism leading to a global well-posedness theory of master equations, beyond [MZ24], and the main results from this reference are a consequence of this general theory.
@@
    $H_{\varphi}(x,\mu,p):=H(x,\mu,x-\nabla\varphi(x))+\frac{\beta^{2}+\beta_{0}^{2}}{2}\Delta\varphi(x)$
    and
    $G_{\varphi}(x,\mu):=G(x,\mu)+\varphi(x).$
@@
    We remark that in general Hamiltonians are only defined up to an additive constant. In classical mechanics, this is saying that we may pick any value to correspond to the ‘zero energy’. In the presence of noise the attentive reader will notice that our $H_{\alpha}$ is not the same as the $H_{\varphi}$ defined above, when $\varphi(x)$ is taken to be $\frac{\alpha}{2}|x|^{2}$ . However, this is not an issue as the difference between the two is a constant. In particular, the two Hamiltonians are equivalent. Thus, we could have defined our $H_{\alpha}$ as $H_{\alpha}(x,\mu,p):=H(x,\mu,p-\alpha x)+{\frac{(\beta^{2}+\beta_{0}^{2})d}{2}}\alpha$ which would then be the exact same as $H_{\varphi}$ defined above, however this would introduce unnecessary notational clutter.
@@
2. Preliminaries
@@
    $\int_{\mathbb{R}^{d}}\langle\partial_{xx}G(x,\mu)\xi(x),\xi(x)\rangle{\mathrm{d}}\mu(x)$
    $+\int_{\mathbb{R}^{d}\times\mathbb{R}^{d}}\langle\partial_{x\mu}G(x,\mu,\tilde{x})\xi(x),\xi(\tilde{x})\rangle{\mathrm{d}}\mu(x){\mathrm{d}}\mu(\tilde{x})\geq 0,$
@@
[…]
    $\leq 0,$
@@
$\lambda_{0}\int_{\mathbb{R}^{d}}\langle\partial_{xx}G(x,\mu)\xi(x),\xi(x)\rangle{\mathrm{d}}\mu(x)$
[…]
@@
3. New Well-Posedness Theories for MFG and master equations
@@
    Assumptions 1 and 2 from above are the standing assumptions imposed in [BMM25].
Let us now restate our crucial observation from the introduction in form of a theorem.
Theorem 3.2.
Proof.
@@
[…]
$+2\alpha\partial_{pp}H(x,\mu,p(x))v(x)\bigg]\bigg|^{2}\bigg\}{\mathrm{d}}\mu(x)$
@@
$\partial_{xx}\tilde{H}(x,\mu,p)$ $=\partial_{xx}H(x,\mu,p-\alpha x)-2\alpha\operatorname{Re}(\partial_{xp}H(x,\mu,p-\alpha x))$
$+\alpha^{2}\partial_{pp}H(x,\mu,p-\alpha x),$
@@
[…]
$+\int_{\mathbb{R}^{d}}\big[\partial_{xx}H(x,\mu,p(x))v(x)-2\alpha\partial_{xp}H(x,\mu,p(x))v(x)$
$+\alpha^{2}\partial_{pp}H(x,\mu,p(x))v(x)\big]\cdot v(x){\mathrm{d}}\mu(x)$
$\leq 0,$
@@
[…]
@@
Suppose that $H:\mathbb{R}^{d}\times{\mathscr{P}}_{2}(\mathbb{R}^{d})\times\mathbb{R}^{d}\to\mathbb{R}$ satisfies
$\partial_{pp}H(x,\mu,p)\geq c_{0}^{-1}I,$
for some $c_{0}>0$ and for all $(x,\mu,p)\in\mathbb{R}^{d}\times{\mathscr{P}}_{2}(\mathbb{R}^{d})\times\mathbb{R}^{d}$. Suppose that $\underline{\kappa}(\partial_{xp}H),\left|\partial_{pp}H\right|,\left|\partial_{xx}H\right|,\left|\partial_{p\mu}H\right|$ and $\left|\partial_{x\mu}H\right|$ are finite. Define
@@
For $\alpha\in\left[\alpha^{H}_{-},\alpha^{H}_{+}\right]$, $\mu\in{\mathscr{P}}_{2}(\mathbb{R}^{d})$, $p\in C(\mathbb{R}^{d};\mathbb{R}^{d})$ and for $v\in L^{2}_{\mu}(\mathbb{R}^{d};\mathbb{R}^{d})$ normalized, i.e. $\int_{\mathbb{R}^{d}}|v(x)|^{2}d\mu=1$, we compute
@@
$+\int_{\mathbb{R}^{d}}\big[\partial_{xx}H(x,\mu,p(x))v(x)-2\alpha\partial_{xp}H(x,\mu,p(x))v(x)$
$+\alpha^{2}\partial_{pp}H(x,\mu,p(x))v(x)\big]\cdot v(x){\mathrm{d}}\mu(x)$
@@
[…]
@@
3.1. Our results and previous results on the master equation involving displacement semi-monotone data